OpenStudy (anonymous):

What is the approximate volume of a slice perpendicular to the x-axis? Consider the region bounded by x2 + y2 = 64 with x ≥ 0, and y ≥ 0 . A solid is created so that the given region is its base and cross-sections perpendicular to the x-axis are squares. Set up a Riemann sum

OpenStudy (anonymous):

I get (8-x)^2(Delta x) but its wrong

OpenStudy (anonymous):

@mathmale can you help me please?

OpenStudy (mathmale):

I'd suggest that this problem would be easier to visualize and to solve if you'd at least make a rough sketch up front. First, you might draw a circle of radius 8, and then draw in a sample "slice" or two perpendicular to the x-axis.

OpenStudy (mathmale):

Looking at that slice, contained within the circle, what's a formula for the length of that slice? alternatively, would you care to defend your own formula (8-x)?

OpenStudy (anonymous):

okay i was solving for y and graphing that not actually graphing a circle like you said

OpenStudy (mathmale):

I overlooked the fact that we are restricted to the first quadrant. A sketch is not mandatory, but would be useful in visualizing what's happening in this problem situation.

OpenStudy (mathmale):

Your proposed method, to solve the equation of the circle for y, will work. Unfortunately, the correct result is not 8-x. Care to try again? Solve x^2 + y^2 = 64 for y^2 and then for y.

OpenStudy (anonymous):

x^2+y^2=64 y^2=64-x^2 y=sqrt(64-x^2) y=8-x ?

OpenStudy (mathmale):

the length of each slice is dependent upon y, and y itself is a function of x. y=sqrt(64-x^2) is fine. But you cannot just take the sqrt of 64 and subtract the sqrt of x^2. Try again.

OpenStudy (anonymous):

so y= sqrt(64-x^2)

OpenStudy (mathmale):

Actually, there's nothing more to do; you already have the correct result, which is y=sqrt(64-x^2). Leave it as is; you cannot reduce or simplify it further. So, what is the area of the individual slice? the thickness? the volume?

OpenStudy (anonymous):

area would be (sqrt(64-x^2))^2 which equals 64-x^2

OpenStudy (anonymous):

volume is (64-x^2)(Delta x)

OpenStudy (mathmale):

That's good. Why don't we check your expression for y by letting x=8? What would y be in that case? What if x were 0? What would y be in that case?

OpenStudy (anonymous):

y = 0 and y = 64

OpenStudy (mathmale):

Check that 64, please. What about that sqrt operator in your formula for y?

OpenStudy (anonymous):

y =8

OpenStudy (mathmale):

Hint: your (correct) formula for y was y=Sqrt(64-x^2).

OpenStudy (anonymous):

meaning the integral is from 0 to 8 right?

OpenStudy (mathmale):

Excuse me. Right. But I was asking you for the lengths of the slices. the shortest is 0; the longest is 8. Now, have you done Riemann Sum problems before?

OpenStudy (anonymous):

yes but only worked on them briefly

OpenStudy (mathmale):

Wish we had time to start with an easier problem. Let me ask you, at this point: What is the integrand in this case (that is, what is the quantity to be integrated)?

OpenStudy (mathmale):